Each side of the triangle has the same length. The perimeter of the triangle is how many units?
-shows a triangle with a side of 2x+7 and a side with 3x+4 and the other side is blank.

Answer by Theo(13342)

since the lengths of the sides of the triangle are all equal, then you get:

2x+7 = 3x+4

solve for x as follows:

start with 2x+7 = 3x+4

subtract 2x from both sides of the equation and subtract 4 from both sides of the equation to get:

7-4 = 3x-2x

simplify to get 3 = x

the value of x is equal to 3.

each side of the triangle will be equal to 2x+7 = 6+7 = 13.

you could have used the other formula as well because 3x+4 = 9+4 = 13.

since all the sides are equal, the third side is also equal to 13.